OpenStudy (anonymous):

find the perimeter of the triangle with the given vertices. (-5,-2),(-1,4),(3,-1) round your solution to two decimal places.

OpenStudy (anonymous):

use the distance formula to find each side ...squareroot of (xsub2-xsub1)+(ysub2-ysub1)

OpenStudy (amistre64):

(-5,-2),(-1,4),(3,-1) 5 2 5 2 5 2 ------------------ 0,0 4,6 8,1 (-5,-2),(-1,4),(3,-1) 1 -4 1-4 1 -4 ------------------ -4,-6 0,0 4,-5 (-5,-2),(-1,4),(3,-1) -3 1 -3 1 -3 1 ----------------- -8,1 -4,5 0,0 \[P=\sqrt{16+36}+\sqrt{64+1}+\sqrt{16+25}\]

OpenStudy (amistre64):

i coulda stopped at twice, but i had to make sure :)

OpenStudy (anonymous):

\[\sqrt{(-5+1)^{2}+(-2-4)^{2}} + \sqrt{(3+1)^{2}+(-1-4)^{2}} + \sqrt{(3+5)^{2}+(-1+2)^{2}}=21.6765\]

OpenStudy (anonymous):

thanks